### v1

@@ Abstract
To ensure the safety of marine and coastal structures, extreme (design) values should be known at the design stage. But for such complex systems, estimating the magnitude of events which are both non-linear and rare is extremely challenging, and involves considerable computational cost to capture the high-fidelity physics. To address this challenge, we offer a new multi-fidelity screening method, Probabilistic Adaptive Screening (PAS), which accurately predicts extreme values of strongly non-linear wave-induced loads while minimising the required high-fidelity simulation duration. The method introduces a probabilistic approach to multi-fidelity screening, allowing efficient linear potential flow indicators to be used in the low-fidelity stage, even for strongly non-linear load cases. The method is validated against a range of cases, including non-linear waves, and ship vertical bending moments, green water impact loads, and slamming loads. It can be concluded that PAS accurately estimates both the short-term distributions and extreme values in all test cases, with most probable maximum (MPM) values within 10% of the available full brute-force Monte-Carlo Simulation (MCS) results. In addition, PAS achieves this performance very efficiently, requiring less than 4% of the high-fidelity simulation time needed for conventional MCS. These results demonstrate that PAS can reliably reproduce the statistics of both weakly and strongly non-linear extreme load problems, while significantly reducing the associated computational cost. The present study validates the statistical PAS framework; further work should focus on validating the full procedure including CFD load simulations, and on validating it for long-term extremes.
@@ 1.1 Wave impacts
A wave impact on a ship, offshore, or coastal structure is a dynamic loading event that occurs when a water wave hits the structure, causing a rapid transfer of momentum. It may occur when such a maritime structure encounters large and steep waves, experiences large wave-induced motions, or a combination of both (see Figure˜1). The resulting loads can cause significant damage, endanger crew or inhabitants, cause pollution of the sea by lost cargo, decrease performance of the structure, or (in rare occasions) endanger the structure itself. Examples include green water and slamming on ships, wave-in-deck loads on offshore structures, and impacts on wind turbines, breakwaters, dams, jetties, bridges, and other coastal infrastructure. Severe wave impact accidents are documented on sailing ships [DG2001, K2018, ABCViking, GuardianMaud] and offshore structures [HOELAB2022, EK2000, ZXZC2025, VAOZ2016, Havtil2025]. Wave impacts also affect the structural reliability of various coastal structures, ranging from breakwaters [CLST2011, ZYMCZ2023], offshore wind turbines [HWZHSL2025], bridges [CST2009], port infrastructure [LLTL2024], lighthouses [ABDRBPD2021], to residential buildings [MECRSLK2024, SES2015]. This makes accurate prediction of the distributions and extreme values of strongly non-linear loads, such as those due to wave impacts, essential. Waves and wave-induced processes are stochastic, so designing for them requires probabilistic design methods. The aim of such methods is not to avoid all accidents, but to ensure that the probability of failure of the considered structure remains below an acceptable level.
@@ 1.2 Requirements for accurate prediction of extreme wave impact loads
As also discussed in [VS2025], obtaining extreme (or design) values for wave impact loads is very challenging because such events are both rare and complex. Converged statistics of rare events can only be obtained from long-duration analysis, whereas we need high-fidelity (HF) simulations to resolve complex physics. HF models in this context can be Computational Fluid Dynamics (CFD) or basin experiments. The ability of CFD to reproduce wave impact loads in a given wave event improved rapidly in recent years (see e.g., [BHV2020]), but simulating durations in the order of the lifetime of a maritime structure is not yet feasible in practical design contexts. Using a dedicated extreme value prediction method (EVPM) can be used to reduce the HF simulation duration. Such an EVPM for strongly non-linear loads should:
@@
[t!] $^{\prime} / x′′{x}^{\prime\prime}$ LF / HF version of variable $x$ RWE Relative Wave Elevation $n_{w}$ # wave encounters in MCS $\widehat{x}$ MPM of variable $x$ VBM Vertical Bending Moment $N$ # seeds in MCS $P_{\text{exp}}$ PoE level corr. to $T_{\text{exp}}$ and $T_{p,e}$ AS Adaptive Screening $C$ Wave crest height $R_{19p}$ RWE at station 19 (from potential flow) CFD Computational Fluid Dynamics $C(u,v)$ Fitted copula model $T_{\text{exp}}$ Target / exposure duration EVPM Extreme Value Prediction Method $\left[\mathbf{d}^{*},\mathbf{h}^{*}\right]$ Predicted HF distrib. $T_{p}$ / $T_{p,e}$ Peak wave (encounter) period HF High-Fidelity $\left[\mathbf{d}^{\text{cop}}_{H,b},\mathbf{h}_{b}^{\text{cop}}\right]$ Distrib. HF copula draw $b$ $V_{s}$ Forward speed of the ship LF Low-Fidelity $\left[\mathbf{d}^{\text{mcs}}_{L},\mathbf{l}^{\text{mcs}}\right]$ Distrib. all LF samples in MCS $V_{hog}$ / $V_{sag}$ Hogging / sagging VBM MCS Monte-Carlo Simulation $\left[\mathbf{d}^{\text{sel}}_{L},\mathbf{l}^{\text{sel}}\right]$ Distrib. present LF samples $z$ # bootstrapping copula draws MD Acquisition function $F_{x}$ Green water force $\epsilon_{1}$ / $\epsilon_{2}$ Case-specific stopping criteria limits MVPD Acquisition function $F_{p41}$ Bow-flare slamming force $\mu$ Wave heading w.r.t. structure MPM Most Probable Maximum $\mathbf{h}^{\text{sel}}$ Present HF samples $\omega$ / $\omega_{e}$ Wave (encounter) frequency PAS Probabilistic Adaptive Screening $m$ # present HF samples Most important nomenclature, where distrib. = distribution.
@@ 1.3 Existing methods and their shortcomings
A literature review of maritime EVPMs can be found in [VS2023], and [VS2025] added recent literature on both EVPMs and related reliability methods. Summarising, some existing methods satisfy most EVPM criteria, but their implementations remain case-specific and untested for strongly wave impact loads. Given these gaps, industry practice for extreme wave impact loads follows standards (e.g., [DNV2019, DNVGL-CG-0130, OTG14, BV2015, BV2019, ABS2020, ABS2021, ITTC2017Ev]), specifying plate thicknesses, or alternative direct analysis methods using experiments with multiple 0.5 to 3 h wave seeds for short-term extreme response estimates.
@@
To address the limitations of existing methods relative to the requirements in Section˜1.2, we previously developed the Adaptive Screening (AS) method [VS2025]. AS is a multi-fidelity approach for predicting extreme non-linear HF wave-induced loads on maritime structures. It uses an LF variable as an indicator to select critical wave events, which are then evaluated with an HF tool to obtain non-linear loads. The framework constructs a surrogate HF distribution from these events, from which probabilistic design loads are derived. The procedure is iterative, adding new HF samples until convergence. AS was validated through a weakly non-linear case (second-order waves) and a moderately non-linear case (vertical bending moments), and further assessed in a third pilot study on a strongly non-linear case (green water impact loads on a containership) [^1] In the AS papers, we used non-linear responses to emphasise the method’s applicability to hydrodynamic motions, accelerations, green water loads etc. However, as discussed in Section 1.3, extreme value prediction aligns with reliability analysis, where loads is the standard term. We therefore now use loads for all hydrodynamic responses.. Further work [VS2025prads] studied the effect of the acquisition function, one of the elements of AS. Most data and all scripts are available in [VSdata2025, VSdata2025PRADS]). The strength of AS for strongly non-linear loads lies in its flexibility (it allows for a wide range of LF indicator variables, including non-linear ones) and in its efficiency (it is optimised to minimise the number of HF simulations). These make the method well-suited for problems with significant non-linearity and sparse datasets. This was supported by [VS2025], which demonstrated good performance for a range of non-linear cases. However, the method also has limitations for complex problems. Selecting an appropriate indicator may require a relatively expensive ‘LF’ model, such as the coarse-mesh CFD used in Case 3 of [VS2025]. This produced accurate results, but the LF MCS became costly. While AS remains far more efficient than a full HF MCS, this additional cost can reduce its practicality in early design stages when computational resources are limited. A poorly chosen indicator may also lead to a non-conservative bias in the estimated design loads. This can be mitigated with a safety factor, but this is not ideal. Thus, AS meets all requirements in Section˜1.2 for simple cases, but it cannot always be efficient and accurate simultaneously in complex cases (it either fails on accuracy or on efficiency).
@@
The existing methods (including AS) generally do not account for the full dependence structure and joint extremes of the LF and HF variables. This may be the reason for some of their shortcomings, for instance the bias introduced by AS when it is applied to complex cases. Copula models offer a flexible way to describe such dependencies. A copula model is a statistical framework for multivariate distributions that decouples marginal behavior from dependence. Compared to directly fitting a joint distribution, copulas can more effectively capture tail dependence, even with limited samples. We included the basics of copula theory in Appendix˜B. Copula modelling of joint dependencies was e.g., recently used in reliability analysis of slopes in spatially variable soils based on limited sample data ([XL2026, WWZCJ2025]), in reliability analysis of systems consisting of dependent equipment parts ([QXSH2025, BCLJ2025]), in reliability-based design optimisation for correlated variables ([ZZYBW2026, LFFT2025]) and to study the joint risk of different relevant variables for the overtopping of breakwaters ([MVM2024]). In extreme value analysis, copulas were used in similar ways to model the joint tail distribution of different variables (e.g., [KJ2019]) and to model the dependence of loads in a mooring line and sea state parameters ([SACM2025]). Copula models have also been used by oceanographers to model the joint statistics of dependent environmental variables (e.g., [VFDKCG2024, DWLSWZ2024, FXLL2024, DV2005]) or to handle spatial dependency between extreme environmental variables (e.g., [BGJM2020, GM2010, DPR2012]). Considering wave impact loading, copula models were also used to model the joint distribution of the quasi-static and impulsive contributions of experimentally measured vertical wave impact pressures on a deck box [WZX2022], to model wave impact maxima on a coastal bridge conditioned on a minimum wave impact rise time [SC2011], to model extreme wave-induced stresses simultaneously measured at two different deck locations of a container vessel [GSN2016], to model the interaction of different parameters of an experimental dam break wave impact event [SL2024], to derive the most likely wave impact load profile on a floating wind turbine as a function of the wave parameters [YXSZB2025], and to model the joint probability of the maxima and rise times of wave impact loads on breakwaters [CPA2011]. Copula models have also been combined with multi-fidelity modelling before: [KNPVW2019, PCMW2016] defined multi-fidelity forms of importance sampling, and [Proppe2019] used multi-fidelity modelling for reliability problems combined with a copula model to establish the model hierarchy. Earlier studies also combined copula modelling with adaptive sampling: [LNDH2025, MZCLGG2026] combined Bayesian model updating with Monte Carlo Simulation (MCS) and copulas in reliability analysis, and [BHRM2018] used a Bayesian model with a copula to capture the joint annual extremes of climate variables that change over time.
@@
Despite their promise, copula-based approaches have, to our knowledge, not yet been combined with multi-fidelity or adaptive sampling frameworks to improve the efficiency of extreme value prediction (or reliability analysis) for wave impact loading (or other strongly non-linear loads). The work done by [KNPVW2019, PCMW2016] is closest to this, but these studies rely on importance sampling rather than adaptive sampling and have not been tested on real-world strongly non-linear cases similar to the wave impact problem. Adaptive sampling is generally considered more flexible and robust than importance sampling for complex tail behaviour.
@@ 1.4 Paper objectives & contribution
As discussed above and in [VS2025], we need a new EVPM for wave impact loads. Serious incidents with maritime structures still occur at sea, and the existing EVPMs are not suitable and / or validated for this problem. We propose a modification to AS that addresses its limitations in complex, strongly non-linear cases. AS relies on LF order statistics as surrogates for HF order statistics, making performance highly sensitive to the choice of LF indicators (often requiring relatively expensive ‘LF’ models). Our new approach keeps the efficient screening and adaptive sampling components, but replaces the order-statistics assumption with a copula model linking LF indicators and HF loads. This enables us to derive the full HF distribution from a limited HF sample, reducing sensitivity to indicator choice, decreasing bias, and ensuring any remaining bias is conservative (unlike the non-conservative AS approach). Consequently, the modification reduces dependence on relatively expensive LF models (e.g., allowing the use of indicators from linear potential flow instead of from coarse mesh CFD), while improving both accuracy and applicability in design contexts. This especially benefits strongly non-linear problems, where suitable LF indicators for AS are computationally costly to obtain. The related objectives of this paper are to:
@@ 1.5 Layout
Section˜2 presents the new PAS method, and Section˜3 introduces four case studies that are used to validate PAS, and to compare it to AS. These include two applications from [VS2025] (second-order waves and vertical bending moments), and add two new strongly non-linear wave impact cases (green water impact forces and bow-flare slamming forces on a sailing ship). Subsequently, Section˜4 explains the method and metrics we used to validate the method’s performance, after which Section˜5 presents and discusses the results. Finally, Section˜6 draws conclusions.

### v2

@@ Abstract
Wave impact loads on a maritime structure can cause casualties, damage, pollution of the sea and operational delays. Their extreme (or design) values should therefore be considered in the design of these structures. However, this is challenging because these events are both rare and complex, requiring high-fidelity computation and long analysis durations to obtain such design loads. Existing extreme value prediction methods are not tailored or validated for wave impacts. We therefore introduce the new Probabilistic Adaptive Screening (PAS) method for predicting extreme non-linear loads on maritime structures. The method introduces a probabilistic approach to multi-fidelity screening, allowing efficient linear potential flow indicators to be used in the low-fidelity stage, even for strongly non-linear load cases. The method is validated against a range of cases, including non-linear waves, ship vertical bending moments, green water impact loads, and slamming loads. It can be concluded that PAS accurately estimates both the short-term distributions and extreme values in these test cases, with most probable maximum (MPM) values within 10% of the available full brute-force Monte-Carlo Simulation (MCS) results. In addition, PAS achieves this performance very efficiently, requiring less than 4% of the high-fidelity simulation time needed for conventional MCS. These results demonstrate that PAS can reliably reproduce the statistics of both weakly and strongly non-linear extreme load problems, while significantly reducing the associated computational cost. The present study validates the statistical PAS framework; further work should focus on validating the full procedure including CFD load simulations, and on validating it for long-term extremes.
@@ 1.1 Wave impacts
A wave impact on a ship, offshore, or coastal structure is a dynamic loading event that occurs when a water wave hits the structure, causing a rapid transfer of momentum. It may occur when such a maritime structure encounters large and steep waves, experiences large wave-induced motions, or some combination of both (see Figure˜1). The resulting loads can cause significant damage, endanger crew or inhabitants, cause pollution of the sea by lost cargo, decrease performance of the structure, or (in rare occasions) endanger the structure itself. Examples include green water and slamming on ships, wave-in-deck loads on offshore structures, and impacts on wind turbines, breakwaters, dams, jetties, bridges, and other coastal infrastructure. Severe wave impact accidents are documented on sailing ships [DG2001, K2018, ABCViking, GuardianMaud] and offshore structures [HOELAB2022, EK2000, ZXZC2025, VAOZ2016, Havtil2025]. Wave impacts also affect the structural reliability of various coastal structures, ranging from breakwaters [CLST2011, ZYMCZ2023], offshore wind turbines [HWZHSL2025], bridges [CST2009], port infrastructure [LLTL2024], lighthouses [ABDRBPD2021], to residential buildings [MECRSLK2024, SES2015]. This makes accurate prediction of the distributions and extreme values of strongly non-linear loads, such as those due to wave impacts, essential. Waves and wave-induced processes are stochastic, so designing for them requires probabilistic design methods. The aim of such methods is not to avoid all accidents, but to ensure that the probability of failure of the considered structure remains below an acceptable level.
@@ 1.2 Requirements for accurate prediction of extreme wave impact loads
As also discussed in [VS2025], obtaining extreme (or design) values for wave impact loads is very challenging because such events are both rare and complex. Converged statistics of rare events can only be obtained from long-duration analysis, whereas we need high-fidelity (HF) simulations to resolve complex physics. HF models in this context can be Computational Fluid Dynamics (CFD) or basin experiments. The ability of CFD to reproduce wave impact loads in a given wave event improved rapidly in recent years (see e.g., [BHV2020]), but simulating durations in the order of the lifetime of a maritime structure is not yet feasible in practical design contexts. Using a dedicated extreme value prediction method (EVPM) can reduce the HF simulation duration, and such an EVPM for strongly non-linear loads should:
@@
[t!] $^{\prime} / x′′{x}^{\prime\prime}$ LF / HF version of variable $x$ RWE Relative Wave Elevation $n_{w}$ # wave encounters in MCS $\widehat{x}$ MPM of variable $x$ VBM Vertical Bending Moment $N$ # seeds in MCS $P_{\text{exp}}$ PoE level corr. to $T_{\text{exp}}$ and $T_{p,e}$ AS Adaptive Screening method $C$ Wave crest height $R_{19p}$ RWE at station 19 (from potential flow) CFD Computational Fluid Dynamics $C(u,v)$ Fitted copula model $T_{\text{exp}}$ Target / exposure duration EVPM Extreme Value Prediction Method $\left[\mathbf{d}^{*},\mathbf{h}^{*}\right]$ Predicted HF distrib. $T_{p}$ / $T_{p,e}$ Peak wave (encounter) period HF High-Fidelity $\left[\mathbf{d}^{\text{cop}}_{H,b},\mathbf{h}_{b}^{\text{cop}}\right]$ Distrib. HF copula draw $b$ $V_{s}$ Forward speed of the ship LF Low-Fidelity $\left[\mathbf{d}^{\text{mcs}}_{L},\mathbf{l}^{\text{mcs}}\right]$ Distrib. all LF samples in MCS $V_{hog}$ / $V_{sag}$ Hogging / sagging VBM MCS Monte-Carlo Simulation $\left[\mathbf{d}^{\text{sel}}_{L},\mathbf{l}^{\text{sel}}\right]$ Distrib. present LF samples $z$ # bootstrapping copula draws MD Acquisition function $F_{x}$ Green water force $\epsilon_{1}$ / $\epsilon_{2}$ Case-specific stopping criteria limits MVPD Acquisition function $F_{p41}$ Bow-flare slamming force $\mu$ Wave heading w.r.t. structure MPM Most Probable Maximum $\mathbf{h}^{\text{sel}}$ Present HF samples $\omega$ / $\omega_{e}$ Wave (encounter) frequency PAS Probabilistic Adaptive Screening $m$ # present HF samples PoE Probability of Exceedance $n$ # LF indicator peaks in MCS Most important nomenclature, where distrib. = distribution.
@@ 1.3 Existing methods and their shortcomings
A literature review of maritime EVPMs can be found in [VS2023], and [VS2025] added recent literature on both EVPMs and related reliability methods. Summarising, some existing methods satisfy multiple EVPM criteria, but their implementations remain case-specific and untested for strongly wave impact loads. Given these gaps, industry practice for extreme wave impact loads follows standards (e.g., [DNV2019, DNVGL-CG-0130, OTG14, BV2015, BV2019, ABS2020, ABS2021, ITTC2017Ev]), specifying plate thicknesses, or alternative direct analysis methods using experiments with multiple 0.5 to 3 h wave seeds for short-term extreme response estimates.
@@
To address the limitations of existing methods relative to the requirements in Section˜1.2, we previously developed the Adaptive Screening (AS) method [VS2025]. AS is a multi-fidelity approach for predicting extreme non-linear HF wave-induced loads on maritime structures. It uses an LF variable as an indicator to select critical wave events, which are then evaluated with an HF tool to obtain non-linear loads. The framework constructs a surrogate HF distribution from these events, from which probabilistic design loads are derived. The procedure is iterative, adding new HF samples until convergence. AS was validated through a weakly non-linear case (second-order waves) and a moderately non-linear case (vertical bending moments), and further assessed in a third pilot study on a strongly non-linear case (green water impact loads on a containership) [^1] In the AS papers, we used non-linear responses to emphasise the method’s applicability to hydrodynamic motions, accelerations, green water loads etc. However, as discussed in Section 1.3, extreme value prediction aligns with reliability analysis, where loads is the standard term. We therefore now use loads for all hydrodynamic responses.. Further work [VS2025prads] studied the effect of the acquisition function, one of the elements of AS. Most data and all scripts are available in [VSdata2025, VSdata2025PRADS]. The strength of AS for strongly non-linear loads lies in its flexibility (it allows for a wide range of LF indicator variables, including non-linear ones) and in its efficiency (it is optimised to minimise the number of HF simulations). These make the method well-suited for problems with significant non-linearity and sparse datasets. This was supported by [VS2025], which demonstrated good performance for a range of non-linear cases. However, the method also has limitations for complex problems. Selecting an appropriate indicator may require a relatively expensive ‘LF’ model, such as the coarse-mesh CFD used in Case 3 of [VS2025]. This produced accurate results, but the LF MCS became costly. While AS remains far more efficient than a full HF MCS, this additional cost can reduce its practicality in early design stages when computational resources are limited. A poorly chosen indicator may also lead to a non-conservative bias in the estimated design loads. This can be mitigated with a safety factor, but this is not ideal. Thus, AS meets all requirements in Section˜1.2 for simple cases, but it cannot always be efficient and accurate simultaneously in complex cases (it either fails on accuracy or on efficiency).
@@
The existing methods (including AS) generally do not account for the full dependence structure and joint extremes of the LF and HF variables. This may be the reason for some of their shortcomings, for instance the bias introduced by AS when it is applied to complex cases. Copula models offer a flexible way to describe such dependencies. A copula model is a statistical framework for multivariate distributions that decouples marginal behavior from dependence. Compared to directly fitting a joint distribution, copulas can more effectively capture tail dependence, even with limited samples. Copula modelling of joint dependencies was e.g., recently used in reliability analysis of slopes in spatially variable soils based on limited sample data [XL2026, WWZCJ2025], in reliability analysis of systems consisting of dependent equipment parts [QXSH2025, BCLJ2025], in reliability-based design optimisation for correlated variables [ZZYBW2026, LFFT2025] and to study the joint risk of different relevant variables for the overtopping of breakwaters [MVM2024]. In extreme value analysis, copulas were used in similar ways to model the joint tail distribution of different variables (e.g., [KJ2019]) and to model the dependence of loads in a mooring line and sea state parameters [SACM2025]. Copula models have been used by oceanographers to model the joint statistics of dependent environmental variables (e.g., [VFDKCG2024, DWLSWZ2024, FXLL2024, DV2005]) or to handle spatial dependency between extreme environmental variables (e.g., [BGJM2020, GM2010, DPR2012]). Considering wave impact loading, copula models have also been used to model the joint distribution of the quasi-static and impulsive contributions of experimentally measured vertical wave impact pressures on a deck box [WZX2022], wave impact maxima on a coastal bridge conditioned on a minimum wave impact rise time [SC2011], extreme wave-induced stresses simultaneously measured at two different deck locations of a container vessel [GSN2016], the interaction of different parameters of an experimental dam break wave impact event [SL2024], the joint probability of the maxima and rise times of wave impact loads on breakwaters [CPA2011], and to derive the most likely wave impact load profile on a floating wind turbine as a function of the wave parameters [YXSZB2025]. Copula models have also been combined with multi-fidelity modelling before: [KNPVW2019, PCMW2016] defined multi-fidelity forms of importance sampling, and [Proppe2019] used multi-fidelity modelling for reliability problems combined with a copula model to establish the model hierarchy. Earlier studies also combined copula modelling with adaptive sampling: [LNDH2025, MZCLGG2026] combined Bayesian model updating with Monte Carlo Simulation (MCS) and copulas in reliability analysis, and [BHRM2018] used a Bayesian model with a copula to capture the joint annual extremes of climate variables that change over time.
@@
Despite their promise, copula-based approaches have, to our knowledge, not yet been combined with multi-fidelity or adaptive sampling frameworks to improve the efficiency of extreme value prediction (or reliability analysis) for wave impact loading (or other strongly non-linear loads). The work done by [PCMW2016, KNPVW2019] is closest to this, but these studies rely on importance sampling rather than adaptive sampling and have not been tested on real-world strongly non-linear cases similar to the wave impact problem. Adaptive sampling is generally considered more flexible and robust than importance sampling for complex tail behaviour.
@@ 1.4 Paper objectives & contribution
As discussed above and in [VS2025], we need a new EVPM for wave impact loads. Serious incidents with maritime structures still occur at sea, and existing EVPMs are not suitable and / or validated for this problem. Therefore we propose a modification to AS that addresses its limitations in complex, strongly non-linear cases. AS relies on LF order statistics as surrogates for HF order statistics, making performance highly sensitive to the choice of LF indicators (often requiring relatively expensive ‘LF’ models). Our new approach keeps the efficient screening and adaptive sampling components, but replaces the order-statistics assumption with a copula model linking LF indicators and HF loads. This enables us to derive the full HF distribution from a limited HF sample, reducing sensitivity to indicator choice, decreasing bias, and ensuring any remaining bias is conservative (unlike the non-conservative AS approach). Consequently, the modification reduces dependence on relatively expensive LF models (e.g., allowing the use of indicators from linear potential flow instead of from coarse mesh CFD), while improving both accuracy and applicability in design contexts. This especially benefits strongly non-linear problems, where suitable LF indicators for AS are computationally costly to obtain. The related objectives of this paper are to:
@@ 1.5 Layout
Section˜2 presents the new PAS method, and Section˜3 introduces four case studies that are used to validate PAS, and to compare it to AS. These include two applications from [VS2025] (second-order waves and vertical bending moments), and add two new strongly non-linear wave impact cases (green water impact forces and bow-flare slamming forces on a sailing ship). Subsequently, Section˜4 explains the metrics we used to validate the method’s performance, after which Section˜5 presents and discusses the results. Finally, Section˜6 draws conclusions.